What is the angular acceleration of a ball that starts at rest and increases its angular velocity uniformly to 5 rad/s in 10 s?

1
Expert's answer
2022-01-02T18:24:06-0500

Solution:


ω=ωo​+ϵt ϵ = "\frac{\omega}{t}"="\frac{5}{10}"= 0.5 ("\frac{rad}{s^2}")


Answer:


Angular acceleration of a ball is 0.5 "\frac{rad}{s^2}"